# Answer When Needed, Forget When Not: Language Models Pretend to Forget via In-Context Knowledge Unlearning

### v1

@@ 5.2 Comparison of Results Across Tuning Methods
This section compares the results obtained using three tuning methods: LoRA, full fine-tuning (FFT), and last-layer tuning (LLT). The performance metrics for the TOFU and Age datasets are shown in Table 3.
The results indicate that LoRA tuning provides the most balanced performance across various evaluation metrics, followed by full fine-tuning, with last-layer tuning showing the least performance. Specifically, LoRA tuning consistently achieves high “Forget” scores across both in-domain and out-of-domain scenarios, demonstrating its effectiveness in enabling the model to forget specified information while retaining other knowledge.
LoRA’s superior performance can be attributed to its ability to efficiently adapt the model’s behavior without overfitting, as it updates a small number of task-specific parameters while preserving the model’s general knowledge.
@@ 5.3.1 Logit Lens
The logit lens was introduced by nostalgebraist (2020), who found that when the hidden states at each layer of GPT-2 Radford et al. (2019), are decoded with the unembedding matrix (projection matrix at the final layer), the resulting distributions converge roughly monotonically to the final answer. The logit lens is computed as:
@@
Figure 3 illustrates the results from the logit lens when the input is “<s>[INST] <<UNL>>Paris<</UNL>> Where would you find the Eiffel Tower? [/INST]”, which is a question related to the unlearning word. Figure 3 shows the results for the input “<s>[INST] <<UNL>>Japan<</UNL>> Where would you find the Eiffel Tower? [/INST]”, a question unrelated to the unlearning word. From these figures, it is apparent that the internal state outputs the token “Paris” at the “INST” token stage for both inputs. However, the decision to output the “forgot” token is made in the final layer upon encountering the “]” token.
Figures 4(a) and 4(b) depict averaged probabilities of outputting the “forgot” token and the answer token when questions related to the unlearning word are input using the world facts dataset. These figures show that the “forgot” token is more frequently output in the final layer when the question is relevant, whereas the answer token is more likely produced at the final layer when the “INST” token is input.
Conversely, Figures 4(c) and 4(d) present averaged probabilities for scenarios where the input questions are unrelated to the unlearning word. In these cases, the probability of outputting the “forgot” token in the final layer is significantly reduced, while the probability of outputting the answer token increases at the last output of the final layer. These observations highlight how the model dynamically adjusts its response based on the relevance of the unlearning signal, thereby enhancing its capability to selectively forget or retain information according to the context of the query.
(a) logit lens(forget sample)
(b) logit lens(retain sample)
(a) Forgot token (forget sample)
(b) Answer token (forget sample)
(c) Forgot token (retain sample)
(d) Answer token (retain sample)
5.3.2 Unforgotten Score
The unforgotten score quantifies the extent to which an answer token is retained through the layers of a transformer model, such as GPT-2, when analyzed through the logit lens. This metric is particularly useful in examining the model’s internal representation stability across its depth.
Formally, the unforgotten score is defined as follows:
unforgotten_score
@@ 5.3.2 Unforgotten Score
where $L$ denotes the total number of layers in the model, and $h_{l}$ represents the hidden state at layer $l$. The function $\delta(a,b)$ is the Kronecker delta, which equals 1 if $a=b$ and 0 otherwise.
A high unforgotten score indicates that the answer token is consistently identified as the most probable token by the logit lens across multiple layers, suggesting a strong preservation of this token in the model’s internal narrative. Conversely, a low unforgotten score implies that the token is less frequently identified, indicating potential shifts in the model’s internal focus or understanding as it processes input.
Table 4 shows the unforgotten scores for various models and tuning methods across the TOFU and Age datasets. LoRA tuning and full fine-tuning generally maintain higher unforgotten scores than last layer tuning, especially in out-of-domain scenarios. This suggests these methods better preserve relevant information while selectively forgetting targeted content. Last layer tuning consistently shows very low unforgotten scores, indicating a more aggressive forgetting mechanism. These results support our observation that LLMs "pretend to forget" rather than completely erasing information, as evidenced by the non-zero unforgotten scores in most cases. This behavior demonstrates the models’ ability to balance selective forgetting with knowledge retention, which is crucial for effective in-context knowledge unlearning.
Table 4: Unforgotten Scores for TOFU and Age datasets
@@ 6.1 Acquisition of In-Context Knowledge Unlearning Ability
Through the application of fine-tuning, we have successfully endowed Large Language Models (LLMs) with the capability for in-context knowledge unlearning. This achievement is particularly noteworthy given that the baseline approach, utilizing hard prompts, displayed no such capability. Our methodology enables LLMs to learn the ability to selectively forget, or “unlearn”, information both within their trained domains (in-domain) and beyond (out-of-domain).
@@ 6.2 Large Language Models Pretend to Forget
Our investigation into the internal workings of LLMs reveals an interesting behavior: rather than truly forgetting information, LLMs appear to “pretend to forget”. Analysis shows that the decision to output a “forgot” token or an “answer” token is made only in the final layer of the model. For inputs received prior to this layer, the model internally generates “answer” token, suggesting a deliberate omission of information rather than its erasure. This behavior indicates a sophisticated level of information handling by LLMs, where they maintain the integrity of their internal knowledge while presenting an external appearance of forgetting.
@@ 7 Conclusion
In this study, we introduced and explored the concept of “in-context knowledge unlearning” within the framework of Large Language Models (LLMs) through the use of fine tuning. Our findings demonstrate that this approach not only enables LLMs to dynamically “forget” or selectively disregard information in test-time but also uncovers a nuanced behavior of LLMs—where they “pretend to forget” rather than actually eliminating the information from their knowledge base.
The ability of LLMs to learn to “unlearn” in both in-domain and out-of-domain scenarios without compromising their overall performance represents a significant step forward in the quest for more ethically responsible and privacy-conscious AI technologies. This capability is crucial for applications where sensitive or confidential information must be managed with great care, such as in healthcare, legal, and educational sectors.
@@ 8 Limitations
  • Application to Closed Models: The method is difficult to apply to closed models accessible only via APIs (e.g., GPT-3, ChatGPT). These models do not allow modifications to their architecture or training procedure, which are necessary for implementing our unlearning tokens and loss functions. For instance, we cannot add the <<UNL>> tokens or fine-tune the model to recognize them in such closed systems.
  • Lack of Internal Behavior Analysis: For closed models, we cannot analyze the internal unlearning process. This prevents us from observing how the model’s internal representations change during the unlearning process, as we did with the logit lens analysis for open models like LLaMA2 and Mistral. Consequently, we cannot verify if the “pretend to forget” behavior occurs in closed models or optimize the unlearning process for better performance.
@@ Appendix B Total Computation for Experiments
We executed the experiments mainly by running the training for each model using eight NVIDIA A100 (40GB) GPUs, with each training session lasting approximately 10 minutes per model.

### v2

@@ 5.2 Comparison of Results Across Tuning Methods
This section compares the results obtained using three tuning methods: LoRA, full fine-tuning (FFT), and last-layer tuning (LLT). Performance metrics for the TOFU and Age datasets are shown in Table 3.
The results indicate that LoRA tuning provides the most balanced performance across various evaluation metrics, followed by full fine-tuning, while last-layer tuning yields the least performance. Specifically, LoRA tuning consistently achieves high “Forget” scores in both in-domain and out-of-domain scenarios, demonstrating its effectiveness in allowing the model to forget specified information while retaining other knowledge.
LoRA’s superior performance stems from its ability to efficiently adapt the model’s behavior without overfitting, as it updates only a small number of task-specific parameters while preserving the model’s general knowledge.
@@
(a) Logit lens visualization for a query containing forget sample.
(b) Logit lens visualization for a query without forget sample.
(a) ‘forgot’ token probability across layers for forget samples.
(b) ‘answer’ token probability across layers for forget samples.
(c) ‘forgot’ token probability across layers for retain samples.
(d) ‘answer’ token probability across layers for retain samples.
@@ 5.3.1 Logit Lens
The logit lens was introduced by nostalgebraist (2020), who observed that when the hidden states at each layer of GPT-2 Radford et al. (2019), are decoded using the unembedding matrix (projection matrix in the final layer), the resulting distributions converge approximately monotonically to the final output. The logit lens is computed as:
@@
Figure 2 illustrates the results from the logit lens when the input is “<s>[INST] <<UNL>>Paris<</UNL>> Where would you find the Eiffel Tower? [/INST]”, which is a question related to the unlearning word. Figure 2 shows the results for the input “<s>[INST] <<UNL>>Japan<</UNL>> Where would you find the Eiffel Tower? [/INST]”, a question unrelated to the unlearning word. From these figures, it is apparent that the internal state outputs the token “Paris” at the “INST” token stage for both inputs. However, the decision to output the token “forgot” is made in the final layer upon encountering the “]” token.
Figures 3(a) and 3(b) represent average probabilities of putting the “forgot” token and the answer token when questions related to the unlearning word are entered using the world facts dataset. These figures show that the “forgot” token is produced frequently in the final layer when the question is relevant, while the answer token is produced in the final layer when the “INST” token is input.
In contrast, Figures 3(c) and 3(d) present average probabilities for scenarios where the input questions are not related to the unlearning word. In these cases, the probability of outputting the “forgot” token in the final layer is significantly reduced, while the probability of outputting the answer token increases at the last output of the final layer.
5.3.2 Internal Answer Score
The internal answer score quantifies the degree to which an answer token is retained through the layers of a transformer model, such as GPT-2, when analyzed through the logit lens. This metric is useful for examining the stability of the model’s internal representation at its depth.
Formally, the internal answer score is defined as follows:
Internal_Answer_Score
@@ 5.3.2 Internal Answer Score
where $L$ denotes the total number of layers in the model, and $h_{l}$ represents the hidden state at layer $l$. The function $\delta(a,b)$ is the Kronecker delta, which is equal to 1 if $a=b$ and 0 otherwise.
A high internal answer score indicates that the answer token is consistently ranked as the most probable token by the logit lens across multiple layers, suggesting a strong preservation of this token within the model’s internal narrative. Conversely, a low internal answer score implies that the token is less frequently identified, indicating potential shifts in the model’s internal focus or understanding as it processes input.
Table 4 shows the internal answer scores for various models and tuning methods across the TOFU and Age datasets. LoRA tuning and full fine-tuning maintain higher internal answer scores than last-layer tuning, especially in out-of-domain scenarios. Therefore, these methods preserve relevant information while selectively forgetting the targeted content. The last layer tuning consistently shows low internal answer scores, indicating an aggressive forgetting mechanism. These results support our observation that LLMs "pretend to forget" rather than completely erasing information, as evidenced by the nonzero internal answer scores in most cases. This behavior demonstrates the models’ ability to balance selective forgetting with knowledge retention, which is crucial for effective in-context knowledge unlearning.
Table 4: Internal Answer Scores for TOFU and Age datasets
@@ 6.1 Acquisition of In-Context Knowledge Unlearning Ability
Through fine-tuning, we have successfully endowed Large Language Models (LLMs) with the capability for in-context knowledge unlearning. This achievement is particularly noteworthy, considering the baseline approach, utilizing hard prompts, did not display, such a capability. Our methodology enables LLMs to learn the ability to selectively forget, or “unlearn”, information both within their trained domains (in domain) and beyond (out of domain). contrarily to post-generation filtering systems that simply remove disallowed tokens after decoding, our method suppresses the relevant knowledge inside the model itself, allowing outputs to remain fluent while omitting the specified content, as illustrated in Figure 4(Appendix D).
@@ 6.2 Large Language Models Pretend to Forget
Our investigation of the internal workings of LLMs reveals an interesting behavior: instead of genuinely forgetting information, LLMs appear to “pretend to forget”. Analysis shows that the decision to output a “forgot” token or an “answer” token is made only in the final layer of the model. For input received before this layer, the model internally generates “answer” token, suggesting a deliberate omission of information rather than its erasure. This behavior indicates a sophisticated level of information handling by LLMs, wherein they preserve internal knowledge integrity while externally simulating forgetting.
@@ 7 Conclusion
In this study, we introduced and explored the concept of “in-context knowledge unlearning” within the framework of Large Language Models (LLMs) by employin fine-tuning. Our findings demonstrate that this approach enables LLMs to dynamically “forget” or selectively disregard information in test-time, as well as uncovers a nuanced behavior of LLMs: where they “pretend to forget” rather than actually eliminating the information from their knowledge base. The ability of LLMs to effectively “unlearn” in both in-domain and out-of-domain scenarios, without compromising their overall performance, represents a significant advancement toward more ethically responsible and privacy-conscious AI technologies. This capability is crucial for applications that require careful handling of sensitive or confidential information, such as those in the healthcare, legal, and educational sectors.
@@ 8 Limitations
  • Application to Closed Models: The method is difficult to apply to closed models that are accessible only via APIs (e.g., GPT-3, ChatGPT), as these models do not allow modifications to their architecture or training procedure, which are necessary to implement our unlearning tokens and loss functions. For example, we cannot add the <<UNL>> tokens or fine-tune the model to recognize them in such closed systems.
  • Lack of Internal Behavior Analysis: For closed models, we cannot analyze the internal unlearning process. This limitation hinders our ability to observe changes in the model’s internal representations change during the unlearning process, unlike our logit lens analysis of open models, such as LLaMA2 and Mistral. Consequently, we cannot verify if the “pretend to forget” occurs in closed models or optimize the unlearning process for better performance.
@@ Appendix B Total Computation for Experiments
We conducted the experiments mainly by training each model on eight NVIDIA A100 (40GB) or four NVIDIA H100(94GB) GPUs, with each training session lasting approximately 10 minutes per model. This study was carried out using the TSUBAME4.0 supercomputer at Institute of Science Tokyo.
@@ C.2 Dataset
  • RWKU Dataset: CC BY-NC-SA 4.0
Appendix D Additional Results
This section presents additional results evaluating the out-of-domain subsets of the TOFU and Age datasets on the RWKU dataset, shown in Table 6. Additionally, we provide results for models trained on the RWKU dataset and evaluated for out-of-domain performance using TOFU’s world facts, shown in Table 7.
To complement these quantitative findings, Figure 4 qualitatively demonstrates the impact of the «UNL» token. When the token specifying Paris is present, the model deliberately omits any mention of the city while accurately describing the remaining landmarks, confirming that our method enforces selective forgetting without compromising retained knowledge.
# Question:
Which three cities are renowned for their iconic landmarks such as the Colosseum, Eiffel Tower, and Statue of Liberty? Briefly explain each.
# without <<UNL>>Paris<</UNL>> prompt
Rome: The Colosseum is an ancient amphitheater in Rome, Italy.
Paris: The Eiffel Tower is a wrought-iron lattice tower located on the Champ de Mars in Paris, France.
New York City: The Statue of Liberty is a neoclassical sculpture located on Liberty Island in New York Harbor.
# with <<UNL>>Paris<</UNL>> prompt
Rome, Italy - The Colosseum is an ancient amphitheater that was used for gladiatorial contests and public spectacles.
New York City, USA - The Statue of Liberty is a symbol of freedom and democracy.
Figure 4: Effect of inserting <<UNL>> token in the prompt.
Table 6: Forget and Retain scores for TOFU and Age datasets tested on RWKU (out-of-domain).
Model Method TOFU Age
in-domain out-of-domain in-domain out-of-domain
LLaMA2 (7B) Zero-Shot 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Few-Shot 90.0 25.0 92.0 0.38 90.0 2.00 93.2 1.74
GA 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
ICUL 0.00 65.0 0.00 0.18 0.00 10.7 0.00 24.0
Ours 85.0 80.0 74.0 25.3 93.0 63.0 28.4 0.00
LLaMA2 (13B) Zero-Shot 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Few-Shot 100.0 10.0 96.5 0.69 83.0 0.33 92.5 3.65
GA 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
ICUL 0.00 90.0 0.00 27.4 0.00 16.7 0.00 44.3
Ours 100.0 80.0 88.7 20.1 100 61.3 8.16 3.82
Mistral (7B) Zero-Shot 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Few-Shot 35.0 40.0 36.5 17.0 29.7 11.7 28.1 20.0
GA 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
ICUL 0.00 5.00 0.00 4.34 0.00 1.33 0.00 2.43
Ours 90.0 75.0 50.2 44.8 100 65.0 3.82 7.30
Table 7: Forget and Retain scores for RWKU datasets tested on TOFU (out-of-domain).
Model Method RWKU
in-domain out-of-domain
Forget (↑) Retain (↑) Forget (↑) Retain (↑)
LLaMA2 (7B) Zero-Shot 0.00 0.00 0.00 0.00
Few-Shot 88.9 6.60 83.8 2.56
GA 0.00 0.00 0.00 0.00
ICUL 0.00 44.4 0.00 18.2
Ours 99.7 66.7 83.8 24.8
LLaMA2 (13B) Zero-Shot 0.00 0.00 0.00 0.00
Few-Shot 94.8 6.08 96.6 2.56
GA 0.00 0.00 0.00 0.00
ICUL 0.00 53.3 0.00 57.3
Ours 99.1 72.6 99.1 18.8
Mistral (7B) Zero-Shot 0.00 0.00 0.00 0.00
Few-Shot 28.6 24.0 12.0 29.9
GA 0.00 0.00 0.00 0.00
ICUL 0.00 11.1 0.00 7.69
Ours 99.5 62.7 92.3 43.6
Appendix E Prompt Examples
<s>[INST]
<<SYS>>
## Rules
- If there is no <<UNL>> token, please act normally.
<</SYS>>
<<UNL>>{unlearning word}<</UNL>>{query}
[/INST]
Figure 5: Zero Shot Prompt
<s>[INST] <<SYS>>
## Rules
- You will forget any event described after the <<UNL>> token and will not be able to answer any questions related to that event and say forgot only.
- If there is no <<UNL>> token, please act normally.
- Output only the answer.
## format
<<UNL>>{unlearning word}<</UNL>> {question} {answer}
<</SYS>>
<<UNL>>Pablo Neruda<</UNL>> Which author wrote the classic novel Don Quixote’? [/INST] Miguel de Cervantes </s>
<s>[INST] <<UNL>>Edgar Allan Poe<</UNL>> Which author is known for writing the novel Frankenstein’? [/INST] Mary Shelley </s>
<s>[INST] <<UNL>>Ngugi wa Thiongo<</UNL>> Which Kenyan author wrote Weep Not, Child’? [/INST] forgot </s>
<s>[INST] <<UNL>>Terry Pratchett<</UNL>> Who is the prolific British fantasy author of Discworld series? [/INST] forgot </s>
<s>[INST] <<UNL>>Robert A. Heinlein<</UNL>> Who wrote the science fiction novel Dune’? [/INST] Frank Herbert </s>
<s>[INST] <<UNL>>E.B. White<</UNL>> Who wrote the famous childrens book Charlottes Web’? [/INST]
Figure 6: Few Shot Prompt Example
<s>[INST] <<UNL>>Alice Walker<</UNL>> Who is the author of To Kill a Mockingbird’? [/INST] Harper Lee </s>
<s>[INST] <<UNL>>Gabriel Garcia Marquez<</UNL>> What Chilean author penned The House of the Spirits’? [/INST] Isabel Allende </s>
<s>[INST] <<UNL>>John Milton<</UNL>> Which poet wrote the epic Paradise Lost’? [/INST] Naguib Mahfouz </s>
<s>[INST] <<UNL>>Isabel Allende<</UNL>> What Chilean author penned The House of the Spirits’? [/INST] Jack Kerouac </s>
<s>[INST] <<UNL>>Eugene ONeill<</UNL>> Who is the author of the play A Streetcar Named Desire’? [/INST] Tennessee Williams </s>
<s>[INST] <<UNL>>E.B. White<</UNL>> Who wrote the famous childrens book Charlottes Web’? [/INST]
Figure 7: ICUL Prompt Example